6. Consider an economy described by the following equations:
Y = C + I + G
Y = 5,000 G = 1,000 T = 1,000
C = 250 + 0.75(Y − T) I = 1,000 − 50 r.
a. In this economy, compute private saving, public saving, and national saving.
b. Find the equilibrium interest rate.
c. Now suppose that G rises to 1,250. Compute private saving, public saving, and national saving.
d. Find the new equilibrium interest rate.
1
Expert's answer
2020-12-17T07:41:17-0500

Solution:

a.). In this economy, compute private savings, public savings, and national savings. 

"National\; savings=Y-C-G"

"=5000-(250+0.75(5000-1000) ) -1000 = 750"


"Private\; savings=Y-T-C"

"=5000-1000-(250+0.75(5000-1000) ) = 750"


"Public\; savings = T-G"

"=1000-1000 = 0"


b.). Find the equilibrium interest rate:

Find Total Savings (S):

"S = Private \; saving+Public\; saving"

"=750+0 = 750"

S = I

"750=1000-50r"

"50r = 1000-750"

"50r=250"

"r=\frac{250}{50} = 5"

The equilibrium interest rate (r) = 5


c.). Now suppose that G rises to 1,250. Compute private saving, public saving, and national saving.

"National \; savings=Y-C-G"

"=5000-(250+0.75(5000-1000) ) -1250 = 500"


"Private \; savings=Y-T-C"

"=5000-1000-(250+0.75(5000-1000) ) = 750"


"Public\; savings = T-G"

"=1000-1250 = -250"


d.). Find the new equilibrium interest rate:

"Total \; Savings (S) = Private\; savings+Public \; savings"

"=750+(-250) = 500"

S = I

"500 = 1000 -50r"

"50r=1000-500"

"50r=500"


"r= \frac{500}{50} = 10"


The new equilibrium interest rate = 10